1. what is the pH of a .00300 M HCN solution? Ka of HCN is 6.2 x 10^-10

2. what is the pH of a .100 mL solution containing .50 M NaD and .20 M HF
Answer #1

1)

HCN dissociates as:

HCN -----> H+ + CN-

3*10^-3 0 0

3*10^-3-x x x

Ka = [H+][CN-]/[HCN]

Ka = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((6.2*10^-10)*3*10^-3) = 1.364*10^-6

since c is much greater than x, our assumption is correct

so, x = 1.364*10^-6 M

So, [H+] = x = 1.364*10^-6 M

use:

pH = -log [H+]

= -log (1.364*10^-6)

= 5.8652

Answer: 5.87

Only 1 question at a time please